OpenStudy (anonymous):

How many moles of nickel (Ni) atoms are in 125 g Ni?

OpenStudy (chmvijay):

can you tell Nickel atomic mass

OpenStudy (missmob):

125/58.69 = 2.12 mol 1.2 10^22 atom Hg= 1.22 10^22/(6.023 10^23) mole= 1.22 10^22/(6.023 10^23) x 200.59 gr = 39.29 gr 1 mole = 6.023 10^23 atoms